# Reconciling Kaplan and Chinchilla Scaling Laws

### v2

@@ 7 Discussion
This note aimed to explain the difference between the Kaplan and Chinchilla scaling coefficients. We found two issues in Kaplan’s study that combined to bias their estimated scaling coefficients; their choice to count only non-embedding parameters, and studying smaller sized model sizes. This means there is curvature in the true relationship between $N_{\setminus E}$ $N_{T}$ (Figure 5). At larger values of $N_{T}$, the embedding parameter counts become negligible, $N_{T}=N_{\setminus E}$, and differences would not arise. Alternatively, had Kaplan studied relationships directly in terms of $N_{T}$, this issue would also not arise, even at this smaller scale (confirmed by our Experiment 1 finding $N_{T}\propto C_{T}^{0.49}$ even for $N_{T}<5M$). The form Kaplan used to predict loss from compute further contributed to differences in the reported compute-loss scaling coefficients.
@@
Inconsistency across scaling studies. Existing literature on scaling is not consistent in its use of non-embedding vs. total compute. Some studies [Henighan et al., 2020, Gordon et al., 2021, Ghorbani et al., 2021, Fernandes et al., 2023, Hu et al., 2024, Bi et al., 2024, Su et al., 2024] follow Kaplan’s approach, using non-embedding parameters or compute, while others [Clark et al., 2022, Que et al., 2024, Wang et al., 2023] adhere to the Chinchilla approach, using total parameters. Our work indicates that this choice can substantially alter scaling exponents, complicating cross-study comparisons. Similarly, the choice of compute-loss equation varies through the literature. Studies such as [Clark et al., 2022, Brown et al., 2020, Smith and Doe, 2024] opt for the Kaplan compute-loss form without offsets. In contrast, [Henighan et al., 2020, Gordon et al., 2021, Fernandes et al., 2023, Hu et al., 2024, Wang et al., 2023] employ the Chinchilla compute-loss form with non-zero offsets. Again, our work suggests that these methodological differences can lead to significant variations in scaling predictions and interpretations.
@@
The lack of a standardized approach in scaling studies risks making comparisons misleading and insights less clear. We see our work as helping to understand certain decisions made in previous studies that should be standardized. Concretely, we advise future studies to report total, rather than non-embedding, parameters, and to include an offset in the compute-loss fitting models. We discuss motivation for these choices below. Furthermore, our initial evidence does not support using multiple cosine decays per model size – we find a single fixed learning rate per model size is sufficient.
@@
Why should embedding parameters contribute to scaling behavior? Several works evidence that embedding parameters capture meaningful language properties. Word embeddings can be factorized into semantically interpretable factors (even the shallow Word2vec) [Mikolov et al., 2013a, b, Arora et al., 2018]. LLMs learn linear embeddings of space and time across scales [Gurnee and Tegmark, 2024]. Developing such meaningful embedding structures allows LLMs to perform high-level language operations, such as arithmetic [McLeish et al., 2024]. Therefore, if one believes that the embedding layer does more than just ‘translate’ tokens to a vector of the correct dimension, we see no reason to exclude them in the parameter count.
@@
Why should a non-zero offset be used in loss-compute predictions? The Chinchilla compute-loss form with a non-zero offset (Eq. 41), is a more appropriate form from the perspective of statistical learning. This approach accounts for the concept of irreducible risk [Hernandez et al., 2023], which posits a lower bound on achievable loss regardless of model or dataset size. This may arise from various factors: inherent biases or limitations in the learning algorithm, or noise in the original task. As a concrete example in language modeling, the best a model can do for the prediction of the first token in a sequence, is to estimate the marginal distribution of all tokens, which leads to a non-zero loss.
@@ Appendix A Appendix
A.1 Derivation of compute-loss analytical form
@@
A.2 Compute-loss coefficient derivation

### v3

@@ 7 Discussion
This paper aimed to explain the difference between the Kaplan and Chinchilla scaling coefficients. We found two issues in Kaplan’s study that combined to bias their estimated scaling coefficients; their choice to count only non-embedding parameters, and studying smaller sized model sizes. This means there is curvature in the true relationship between $N_{\setminus E}$ $N_{T}$ (Figure 5). At larger values of $N_{T}$, the embedding parameter counts become negligible, $N_{T}=N_{\setminus E}$, and differences would not arise. Alternatively, had Kaplan studied relationships directly in terms of $N_{T}$, this issue would also not arise, even at this smaller scale (confirmed by our Experiment 1 finding $N_{T}\propto C_{T}^{0.49}$ even for $N_{T}<5M$). The form Kaplan used to predict loss from compute further contributed to differences in the reported compute-loss scaling coefficients.
@@
Inconsistency across scaling studies. Existing literature on scaling is not consistent in its use of non-embedding vs. total compute. Some studies (Henighan et al., 2020; Gordon et al., 2021; Ghorbani et al., 2021; Fernandes et al., 2023; Hu et al., 2024; Bi et al., 2024; Su et al., 2024) follow Kaplan’s approach, using non-embedding parameters or compute, while others (Clark et al., 2022; Que et al., 2024; Wang et al., 2023) adhere to the Chinchilla approach, using total parameters. Our work indicates that this choice can substantially alter scaling exponents, complicating cross-study comparisons. Similarly, the choice of compute-loss equation varies through the literature. Studies such as (Clark et al., 2022; Brown et al., 2020; Smith & Doe, 2024) opt for the Kaplan compute-loss form without offsets. In contrast, (Henighan et al., 2020; Gordon et al., 2021; Fernandes et al., 2023; Hu et al., 2024; Wang et al., 2023) employ the Chinchilla compute-loss form with non-zero offsets. Again, our work suggests that these methodological differences can lead to significant variations in scaling predictions and interpretations.
@@
The lack of a standardized approach in scaling studies risks making comparisons misleading and insights less clear. We see our work as helping to understand certain decisions made in previous studies that should be standardized. Concretely, we advise future studies to report total, rather than non-embedding, parameters, and to include an offset in the compute-loss fitting models. We discuss motivation for these choices below. Furthermore, our initial evidence does not support using multiple cosine decays per model size – we find a single fixed learning rate per model size is sufficient for measuring compute-optimal parameter coefficients.
@@
Why should embedding parameters contribute to scaling behavior? Several works evidence that embedding parameters capture meaningful language properties. Word embeddings can be factorized into semantically interpretable factors (even the shallow Word2vec) (Mikolov et al., 2013a; b; Arora et al., 2018). LLMs learn linear embeddings of space and time across scales (Gurnee & Tegmark, 2024). Developing such meaningful embedding structures allows LLMs to perform high-level language operations, such as arithmetic (McLeish et al., 2024). Therefore, if one believes that the embedding layer does more than just ‘translate’ tokens to a vector of the correct dimension, we see no reason to exclude them in the parameter count.
@@
Why should a non-zero offset be used in loss-compute predictions? The Chinchilla compute-loss form with a non-zero offset (Eq. 43), is a more appropriate form from the perspective of statistical learning. This approach accounts for the concept of irreducible risk (Hernandez et al., 2023), which posits a lower bound on achievable loss regardless of model or dataset size. This may arise from various factors: inherent biases or limitations in the learning algorithm, or noise in the original task. As a concrete example in language modeling, the best a model can do for the prediction of the first token in a sequence, is to estimate the marginal distribution of all tokens, which leads to a non-zero loss.
@@ Appendix A Appendix
A.1 Derivation of Eq.20
This section derives Eq.20
First note that
$\frac{d\log(C_{\setminus E})}{d\log(N^{*}_{\setminus E})}$ $=\frac{d\log(C_{\setminus E})}{dN^{*}_{\setminus E}}\frac{dN^{*}_{\setminus E}}{d\log(N^{*}_{\setminus E})}$ (66)
$=\frac{d\log(C_{\setminus E})}{dN^{*}_{\setminus E}}N^{*}_{\setminus E}$ (67)
Recall the definition of $C_{\setminus E}$ from Eq. 19
$C_{\setminus E}$ $=6N^{*}_{\setminus E}\left(N^{*}_{\setminus E}+\frac{\omega}{3}{N^{*}_{\setminus E}}^{1/3}\right)^{-\frac{1}{\beta}}\left(N^{*}_{\setminus E}+\omega{N^{*}_{\setminus E}}^{1/3}\right)^{\frac{1+\alpha}{\beta}}\left(\frac{\beta}{\alpha}\frac{D_{c}}{N_{c}}\right)^{\frac{1}{\beta}}$ (68)
$\log(C_{\setminus E})$ $=\underbrace{\log(N^{*}_{\setminus E})}_{\text{term 1}}-\underbrace{\frac{1}{\beta}\log\left(N^{*}_{\setminus E}+\frac{\omega}{3}{N^{*}_{\setminus E}}^{1/3}\right)}_{\text{term 2}}+\underbrace{\frac{1+\alpha}{\beta}\log\left(N^{*}_{\setminus E}+\omega{N^{*}_{\setminus E}}^{1/3}\right)}_{\text{term 3}}+\text{const.}$. (69)
where const. does not depend on $N^{*}_{\setminus E}$. We now can take the derivative of each term.
Derivative of term 1.
$\frac{d\log(N^{*}_{\setminus E})}{dN^{*}_{\setminus E}}=\frac{1}{N^{*}_{\setminus E}}$ (70)
Derivative of term 2.
$\frac{1}{\beta}\frac{d\log\left(N^{*}_{\setminus E}+\frac{\omega}{3}{N^{*}_{\setminus E}}^{1/3}\right)}{dN^{*}_{\setminus E}}$ $=\frac{1}{\beta}\frac{d\log\left(N^{*}_{\setminus E}+\frac{\omega}{3}{N^{*}_{\setminus E}}^{1/3}\right)}{dN^{*}_{\setminus E}+\frac{\omega}{3}{N^{*}_{\setminus E}}^{1/3}}\frac{dN^{*}_{\setminus E}+\frac{\omega}{3}{N^{*}_{\setminus E}}^{1/3}}{dN^{*}_{\setminus E}}=\frac{1}{\beta}\frac{1+\frac{\omega}{9}{N^{*}_{\setminus E}}^{-2/3}}{N^{*}_{\setminus E}+\frac{\omega}{3}{N^{*}_{\setminus E}}^{1/3}}$ (71)
Derivative of term 3.
$\frac{\alpha+1}{\beta}\frac{d\log\left(N^{*}_{\setminus E}+\omega{N^{*}_{\setminus E}}^{1/3}\right)}{dN^{*}_{\setminus E}}=\frac{\alpha+1}{\beta}\frac{d\log\left(N^{*}_{\setminus E}+\omega{N^{*}_{\setminus E}}^{1/3}\right)}{dN^{*}_{\setminus E}+\omega{N^{*}_{\setminus E}}^{1/3}}\frac{dN^{*}_{\setminus E}+\omega{N^{*}_{\setminus E}}^{1/3}}{dN^{*}_{\setminus E}}=\frac{\alpha+1}{\beta}\frac{1+\frac{\omega}{3}{N^{*}_{\setminus E}}^{-2/3}}{N^{*}_{\setminus E}+\omega{N^{*}_{\setminus E}}^{1/3}}$ (72)
Then assemble all terms and multiply by $N^{*}_{\setminus E}$ as per Eq. 67.
A.2 Derivation of compute-loss analytical form in Eq. 46
@@
A.3 Compute-loss coefficient derivation